A family drove to a resort at an average speed of 25 mph and later returned over the same road at an average speed of 40 mph. Find the distance to the resort if the total driving time was 13 hours.
Answer by TimothyLamb(4379)

s = d/t
d = st
---
average speed for the whole trip:
s = (25 + 40)/2
s = 32.5 mph
---
total distance driven:
d = 32.5*13
d = 422.5 miles
---
distance to resort:
d/2 = 422.5/2
d/2 = 211.25 miles
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php